OpenStudy (melissa_something):

*Will fan and medal!* Given the equation f(x)=2x+6 find:

OpenStudy (melissa_something):

f(x+h) and f(x+h)-f(x) all over/ h

OpenStudy (melissa_something):

Oh and simplify -___-

OpenStudy (anonymous):

f(x)=2x+6 f(x+h)=2(x+h)+6

OpenStudy (melissa_something):

I don't understand :(

OpenStudy (anonymous):

okay f(x+h) means x put (x+h)

OpenStudy (anonymous):

f(x)=2x+6 f(x+h)=2(x+h)+6

OpenStudy (anonymous):

is that clear

OpenStudy (anonymous):

[f(x+h)-f(x)]/h =[2(x+h)+6-(2x+6)]/h =[2x+2h+6-2x-6]/h =2h/h=2

OpenStudy (melissa_something):

It is wrong, the first one :/

OpenStudy (anonymous):

f(x+h)=2(x+h)+6=2x+2h+6

OpenStudy (anonymous):

this is right

OpenStudy (melissa_something):

2x+2h+6?

OpenStudy (melissa_something):

I got 2x+6+fh but that was wrong too lol

OpenStudy (anonymous):

it should be 2x+6+2h

OpenStudy (anonymous):

the (f) there has no meaning, it should be (2)